One canned juice drink is 30% is orange juice, another is 10% orange juice. How many liters of each should be mixed together in order to get 20 Liters that is 21% orange juice?
Answer by stanbon(75887)

One canned juice drink is 30% is orange juice, another is 10% orange juice. How many liters of each should be mixed together in order to get 20 Liters that is 21% orange juice?
-----
Equation:
orange + orange = orange
0.30x + 0.10(20-x) = 0.21*20
-------
30x + 10*20 - 10x = 21*20
20x = 11*20
x = 11 liters (amt. of 30% solution needed)
20-x = 9 liters (amt. of 10% solution needed)
---------------
Cheers,
Stan H.
--------------